A car has an initial speed of 30m/s, and a mass of 2000kg. If the car is stopped in 30 seconds, what is the power required to stop the car? How much power is required if it stops in 10 seconds?


1
Expert's answer
2021-04-18T19:33:20-0400

The power, needed to stop the car is "P = \frac{A}{t}", where "A" is the work done to stop the car, and "t" is the time needed to do it.

According to energy-work relation, here the absolute value of the work needed to stop the car is just cars' initial kinetic energy "\frac{m v^2}{2}". Hence:

"P_1 = \frac{m v^2}{2 t_1} = 30 k W",

"P_2 = \frac{m v^2}{2 t_2} = 90 k W".